# Convergence Analysis of Natural Gradient Descent for Over-parameterized Physics-Informed Neural Networks

### v1

@@ 7.4. Proof of Theorem 2
Corollary 2 implies that when $m$ is large enough, we have $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq 1$ and then $\|\bm{w}_{r}(k+1)\|_{2}\leq B$. Thus, in induction, we only need to prove that (67) also holds for $t=k+1$, which relies on the recursion formula (26).
Recall that
$\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}+\bm{I}_{1}(k).$
From Corollary 2 and Lemma 5, taking $CM^{2}R<\frac{\lambda_{0}}{4}$ in (24) and $R^{{}^{\prime}}\leq R$ in (68) yields that $\lambda_{min}(\bm{H}(k))\geq\frac{\lambda_{0}}{2}$ and
@@
$\left\|\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}\right\|_{2}^{2}$ (69)
$=\left\|(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left\langle(\bm{I}-\eta\bm{H}(k))\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix},\bm{I}_{1}(k)\right\rangle$
$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}.$
Thus, it remains to bound $\|\bm{I}_{1}(k)\|_{2}$.
@@
$I_{1}^{p}(k)=s_{p}(k+1)-s_{p}(k)-\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle,$
@@
$I_{1}^{n_{1}+j}(k)=h_{j}(k+1)-h_{j}(k)-\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\rangle.$
@@
$s_{p}(k)=\frac{1}{\sqrt{n_{1}}}\left(\frac{1}{\sqrt{m}}\sum\limits_{r=1}^{m}a_{r}\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}x_{p})w_{r0}(k)-a_{r}\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}x_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}-f(x_{p})\right)$
@@
$\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+2}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\bm{x}_{p}$
$\quad-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}(k)\end{pmatrix}].$
@@
$\hat{\chi}_{pr}^{1}(k)=\chi_{pr}^{1}(k+1)-\chi_{pr}^{1}(k)-\langle\frac{\partial\chi_{pr}^{1}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\rangle$
@@
$\hat{\chi}_{pr}^{2}(k)=\chi_{pr}^{2}(k+1)-\chi_{pr}^{2}(k)-\langle\frac{\partial\chi_{pr}^{2}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\rangle.$
@@
The purpose of defining them in this way is to enable us to handle the terms related to the operators $\frac{\partial u}{\partial t}$ and $\Delta u$ separately.
Similar to that for regression problems with ReLU activation function, we first define some events. For $p\in[n_{1}]$, define
@@
and let $S_{p}^{\perp}=[n_{1}]\textbackslash S_{p}$.
In the following, we are going to show that $|\hat{\chi}_{pr}^{1}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$ for every $r\in[m]$ and $|\hat{\chi}_{pr}^{2}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$ for $r\in S_{p}$, $|\hat{\chi}_{pr}^{2}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2})$ for $r\in S_{p}^{\perp}$. Thus, we can prove that $\|\bm{I}_{1}\|_{2}=\mathcal{O}\left(\frac{\sqrt{L(k)}}{\sqrt{m}}\right)$. Then combining with (69) leads to a milder requirement for $\eta$.
For $\hat{\chi}_{pr}^{1}(k)$, we have that
@@
For $r\in S_{p}$, as $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq R,\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R$, we have that $I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{p}\geq 0\}=I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}$, which yields that
@@
$I_{1}^{n_{1}+j}(k)$ $=h_{j}(k+1)-h_{j}(k)-\langle\bm{w}(k+1)-\bm{w}(k),\frac{\partial h_{j}(k)}{\partial\bm{w}}\rangle$
@@
Recall that
$A_{pr}=\{\exists\bm{w}:\|\bm{w}-\bm{w}_{r}\|_{2}\leq R,I\{\bm{w}^{T}\bm{x}_{p}\geq 0\}\neq I\{\bm{w}_{r}^{T}\bm{x}_{p}\geq 0\}\}$
and
$P(A_{p,r})\leq\frac{R}{\sqrt{2\pi}},\ S_{p}=\{r\in[m]:I\{A_{p,r}\}=0\}.$
@@
Then the inequality holds for all $p\in[n_{1}]$ with probability at least $1-n_{1}e^{-mR}$. Thus from (78), we have that for every $p\in[n_{1}]$
@@
Plugging it into (69) yields that
$\left\|\begin{bmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{bmatrix}\right\|_{2}^{2}$
$\leq(1-\frac{\eta\lambda_{0}}{2})^{2}\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2(1-\frac{\eta\lambda_{0}}{2})\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}$
$\leq[(1-\frac{\eta\lambda_{0}}{2})^{2}+C^{2}(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R)^{2}+2C(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R)]\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2}$
$\leq(1-\frac{\eta\lambda_{0}}{2})\left\|\begin{bmatrix}\bm{s}(k)\\ \bm{h}(k)\end{bmatrix}\right\|_{2}^{2},$
where the last inequality requires that
@@
Recall that we also require that $CM^{2}R<\frac{\lambda_{0}}{4}$ and
$R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}<R.$
Finally, with $R=\mathcal{O}(\frac{\lambda_{0}}{B^{4}})$ and Lemma 11 for the upper bound of $L(0)$, $m$ needs to satisfies that
$m=\Omega\left(\frac{B^{12}L(0)}{\lambda_{0}^{4}}\right)=\Omega\left(\frac{d^{12}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$

### v3

@@ 9. Proof of Section 3
Proof Sketch: Assume that Condition 2 holds for $t=0,\cdots,k$, it suffices to demonstrate that Condition 2 also holds for $t=k+1.$
From the recursion formula (39), we have that
$\left\|\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}\right\|_{2}^{2}$ (49)
$=\left\|(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}+\bm{I}_{1}(k)\right\|_{2}^{2}$
$\leq\left\|\bm{I}-\eta\bm{H}(k)\right\|_{2}^{2}\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}+\left\|\bm{I}_{1}(k)\right\|_{2}^{2}+2\left\|\bm{I}-\eta\bm{H}(k)\right\|_{2}\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}\left\|\bm{I}_{1}(k)\right\|_{2},$
where the inequality follows from the Cauchy’s inequality.
Combining Corollary B.2 with Lemma 3.5, we can deduce that when $m$ is large enough, we have $\|\bm{H}(k)-\bm{H}(0)\|_{2}\leq\lambda_{0}/4$. Thus, $\lambda_{min}(\bm{H}(k))\geq\lambda_{0}/2$ and $\bm{I}-\eta\bm{H}(k)$ is positive definite when $\eta=\mathcal{O}(1/\|\bm{H}^{\infty}\|_{2})$. On the other hand, with Corollary B.2, we can derive that $\|\bm{I}_{1}(k)\|_{2}=\mathcal{O}(\eta\sqrt{L}(k)/\sqrt{m})$. Plugging these results into (48), we have
$\left\|\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}\right\|_{2}^{2}$ (50)
$=\left(\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}+\mathcal{O}\left(\frac{\eta^{2}}{m}\right)+\mathcal{O}\left(\frac{\eta}{\sqrt{m}}\right)\right)\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}$
$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2},$
where the last inequality holds when $m$ is large enough.
Now we come to prove Theorem 3.7.
Proof.
Corollary B.2 implies that when $m$ is large enough, we have $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq 1$ and then $\|\bm{w}_{r}(k+1)\|_{2}\leq B$. Thus, in induction, we only need to prove that (46) also holds for $t=k+1$, which relies on the recursion formula (49).
Recall that the recursion formula is
$\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}=(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}+\bm{I}_{1}(k).$
From Corollary B.2 and Lemma 3.5, taking $CM^{2}R<\frac{\lambda_{0}}{4}$ in (8) and $R^{{}^{\prime}}\leq R$ in (47) yields that $\lambda_{min}(\bm{H}(k))\geq\lambda_{min}(\bm{H}(0))-\frac{\lambda_{0}}{4}\geq\frac{\lambda_{0}}{2}$ and
@@
$\left\|\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}\right\|_{2}^{2}$ (51)
$=\left\|(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left\langle(\bm{I}-\eta\bm{H}(k))\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix},\bm{I}_{1}(k)\right\rangle$
$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}.$
Thus, it remains only to bound $\|\bm{I}_{1}(k)\|_{2}$.
@@
$I_{1}^{p}(k)=s_{p}(k+1)-s_{p}(k)-\left\langle\frac{\partial s_{p}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle,$
@@
$I_{1}^{n_{1}+j}(k)=h_{j}(k+1)-h_{j}(k)-\left\langle\frac{\partial h_{j}(k)}{\partial\bm{w}},\bm{w}(k+1)-\bm{w}(k)\right\rangle.$
@@
$s_{p}(k)=\frac{1}{\sqrt{n_{1}}}\left(\frac{1}{\sqrt{m}}\left(\sum\limits_{r=1}^{m}a_{r}\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}x_{p})w_{r0}(k)-a_{r}\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}x_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\right)-f(x_{p})\right)$
@@
$\frac{\partial s_{p}(k)}{\partial\bm{w}_{r}}$ $=\frac{a_{r}}{\sqrt{n_{1}m}}\left[\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})w_{r0}(k)\bm{x}_{p}+\sigma^{{}^{\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}1\\ \bm{0}_{d+2}\end{pmatrix}-\sigma^{{}^{\prime\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\|\bm{w}_{r1}(k)\|_{2}^{2}\bm{x}_{p}\right.$
$\quad\left.-2\sigma^{{}^{\prime\prime}}(\bm{w}_{r}(k)^{T}\bm{x}_{p})\begin{pmatrix}0\\ \bm{w}_{r1}(k)\end{pmatrix}\right].$
@@
$\hat{\chi}_{pr}^{1}(k)=\chi_{pr}^{1}(k+1)-\chi_{pr}^{1}(k)-\left\langle\frac{\partial\chi_{pr}^{1}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\right\rangle$
@@
$\hat{\chi}_{pr}^{2}(k)=\chi_{pr}^{2}(k+1)-\chi_{pr}^{2}(k)-\left\langle\frac{\partial\chi_{pr}^{2}(k)}{\partial\bm{w}_{r}},\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\right\rangle.$
@@
The purpose of defining $\hat{\chi}_{pr}^{1}(k)$ and $\hat{\chi}_{pr}^{1}(k)$ in this way is to enable us to handle the terms related to the operators $\frac{\partial u}{\partial t}$ and $\Delta u$ separately.
We first recall some definitions. For $p\in[n_{1}]$,
@@
and $S_{p}=\{r\in[m]:I\{A_{p,r}=0\}\}$, $S_{p}^{\perp}=[n_{1}]\backslash S_{p}$.
In the following, we are going to show that $|\hat{\chi}_{pr}^{1}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$ for every $r\in[m]$ and $|\hat{\chi}_{pr}^{2}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2}^{2})$ for $r\in S_{p}$, $|\hat{\chi}_{pr}^{2}(k)|=\mathcal{O}(\|\bm{w}_{r}(k+1)-\bm{w}_{r}(k)\|_{2})$ for $r\in S_{p}^{\perp}$. Thus, we can prove that $\|\bm{I}_{1}(k)\|_{2}=\mathcal{O}\left(\frac{\sqrt{L(k)}}{\sqrt{m}}\right)$. Then combining with (69) leads to the conclusion.
For $\hat{\chi}_{pr}^{1}(k)$, from its definition, we have that
@@
For $r\in S_{p}$, since $\|\bm{w}_{r}(k+1)-\bm{w}_{r}(0)\|_{2}\leq R,\|\bm{w}_{r}(k)-\bm{w}_{r}(0)\|_{2}\leq R$, we have that $I\{\bm{w}_{r}(k+1)^{T}\bm{x}_{p}\geq 0\}=I\{\bm{w}_{r}(k)^{T}\bm{x}_{p}\geq 0\}$, which yields that
@@
$I_{1}^{n_{1}+j}(k)$ $=h_{j}(k+1)-h_{j}(k)-\left\langle\bm{w}(k+1)-\bm{w}(k),\frac{\partial h_{j}(k)}{\partial\bm{w}}\right\rangle$
@@
Note that
$P(A_{p,r})\leq\frac{2R}{\sqrt{2\pi}},\ S_{p}=\{r\in[m]:I\{A_{p,r}\}=0\}.$
@@
Then the inequality holds for all $p\in[n_{1}]$ with probability at least $1-n_{1}e^{-mR}$. Plugging this into (59), we can conclude that for every $p\in[n_{1}]$
@@
Plugging this into (50) yields that
$\left\|\begin{pmatrix}\bm{s}(k+1)\\ \bm{h}(k+1)\end{pmatrix}\right\|_{2}^{2}$
$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}+\|\bm{I}_{1}(k)\|_{2}^{2}+2\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}\|\bm{I}_{1}(k)\|_{2}$
$\leq\left[\left(1-\frac{\eta\lambda_{0}}{2}\right)^{2}+C^{2}\left(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R\right)^{2}+2C\left(\frac{\eta^{2}B^{5}\sqrt{L(0)}}{\sqrt{m}}+\eta B^{4}R\right)\right]\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2}$
$\leq\left(1-\frac{\eta\lambda_{0}}{2}\right)\left\|\begin{pmatrix}\bm{s}(k)\\ \bm{h}(k)\end{pmatrix}\right\|_{2}^{2},$
where $C$ is a universal constant and the last inequality requires that
@@
Recall that we also require $CM^{2}R<\frac{\lambda_{0}}{4}$ for $R$ in (8) and
$R^{{}^{\prime}}=\frac{CB^{2}\sqrt{L(0)}}{\sqrt{m}\lambda_{0}}<R$
for $R^{{}^{\prime}}$ in (47) to make sure $\|\bm{H}(k)-\bm{H}(0)\|_{2}\leq\frac{\lambda_{0}}{4}$.
Finally, with $R=\mathcal{O}(\frac{\lambda_{0}}{M^{2}})$ and the upper bound of $L(0)$, $m$ needs to satisfies that
$m=\Omega\left(\frac{M^{4}B^{4}L(0)}{\lambda_{0}^{4}}\right)=\Omega\left(\frac{d^{8}}{\lambda_{0}^{4}}\log^{6}\left(\frac{md}{\delta}\right)\log\left(\frac{n_{1}+n_{2}}{\delta}\right)\right).$